1.     How many molecules are in are in 3.5 mol of Ca3(PO4)2? How many oxygen atoms are there?    

 



1
Expert's answer
2021-01-19T04:40:00-0500

Solution:

(1):

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of Ca3(PO4)2 molecules = 3.5 mol × (6.022×1023 molecules / 1 mol) = 2.1077×1024 molecules

Number of Ca3(PO4)2 molecules = 2.11×1024 molecules


(2):

The schematic equation:

Ca3(PO4)2 → 3Ca + 2P + 8O

According to the equation above: n(Ca3(PO4)2) = n(O)/8

n(O) = n(Ca3(PO4)2) × 8 = 3.5 mol × 8 = 28 mol

n(O) = 28 mol


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of O atoms = 28 mol × (6.022×1023 atoms / 1 mol) = 1.6862×1025 atoms

Number of O atoms = 1.69×1025 atoms


Answer:

(1): 2.11×1024 Ca3(PO4)2 molecules are in 3.5 mol of Ca3(PO4)2.

(2): 1.69×1025 oxygen (O) atoms (or 28 mol) are in 3.5 mol of Ca3(PO4)2.